OpenStudy (anonymous):

final exam today. I cannot remember how to answer these kinds of questions! Someone help remind me? Bubba can mow the lawn in 3 hours. Pierre takes 10 hours. How long does it take them to mow the lawn together?

OpenStudy (anonymous):

save time on the exam, go right to the answer \[\frac{3\times 10}{3+10}\] then compute it takes less than a minute

OpenStudy (anonymous):

we can work out the long involved process for solving this, but this is the answer you will get only thing left to do is arithmetic \[\frac{3\times 10}{3+10}=\frac{30}{13}\]

OpenStudy (lgbasallote):

or if you want the formula.. \[\huge \frac tx + \frac ty = 1\] where t is the total time working together x is the time it takes for an individual to finish the job alone y is the time it takes for another individual to finish the job alone

OpenStudy (anonymous):

leaving you with more times for other exam questions

OpenStudy (anonymous):

\[\huge (\frac 1x + \frac 1y)t = 1\] \[\huge(\frac{x+y}{xy})t=1\] \[\huge t=\frac{xy}{x+y}\] as promised

OpenStudy (anonymous):

you are amazing! thank you to both of you!